# Safe and Optimal Learning from Preferences via Weighted Temporal Logic with Applications in Robotics and Formula 1

### v1

@@ V-B Learning-to-rank
When we take a closer look at the learned weights, we see that, for the DNF/DNS excluded case, the most important part of the formula is the initial grid position, followed by lap times, pit times, delta-to-lead, and completion. On the other hand, when DNF/DNS is included, the importance order changes to first the lap times, followed by the completion, initial grid position and pit times, and finally, delta-to-lead. These insights can help teams improve their race-day strategies.
@@
TABLE I: Performance statistics of RS, and the performance of the proposed method on training and test sets.
Accuracy DNF/DNS excluded DNF/DNS included
(%) Train Test Train Test
RS[10] $83.8\%\pm 0.6\%$ $84.8\%\pm 1.3\%$ $78.5\%\pm 0.7\%$ $76.6\%\pm 9.7\%$
Ours $86.87\%$ $84.97\%$ $84.21\%$ $81.05\%$
@@
Finally, we want to answer the following question: given only the laps observed so far until lap $K$, how accurately can we predict the final standings? Figure 3 shows the evolution of the prediction accuracy over the race. More specifically, when DNF/DNS are excluded, as Figure 3(a) shows, the model reaches over $85\%$ accuracy after the first five laps. This observation may indicate that when unpredictable events are not considered, the performance patterns emerge early. On the other hand, when DNF/DNS are included, the accuracy improves more slowly over laps, reflecting the unpredictable nature of the race events.
@@
It is worth emphasizing that the goal of this experiment is not to provide insights on Formula 1 but to serve as a proof-of-concept. The results highlight the abilities of temporal logic learning in analyzing and learning patterns in an interpretable and efficient manner for complex problems like learning-to-rank on real-world sequential data, such as racing.
@@ VI Conclusion, Limitations and Future Work
In this work, we proposed a safe learning-from-human-feedback framework using temporal logic formalism that can be efficiently solved to optimality. The framework consists of two key procedures, structural pruning and $\log$-transform. Structural pruning systematically removes parts of the robustness tree that do not affect the robustness value, reducing the problem size. The $\log$-transform linearizes the robustness computation constraints with respect to weights, enabling an MILP formulation. We demonstrated the performance of the proposed method on two experiments: the preference learning problem for a robot navigation task and the learning-to-rank problem using Formula 1 data. Results show that the method can capture and reflect diverse preferences effectively.
Despite the success of the proposed method in the experiments, several factors may limit its applicability. First, the method requires domain knowledge and expertise in writing temporal logic formulas to define task specifications. While this need introduces a reliance on expert input, it comes with the benefit of interpretability. Second, MILP formulation requires linear objective functions, which may restrict the applicability to problems that involve non-linear objectives or require certain forms of regularization. Looking forward, we plan to integrate more streamlined task specification methods, potentially using large language models to translate plain sentences to STL formulas.

### v2

@@ V-B Learning to Rank
When we take a closer look at the learned weights, we see that, for the DNF/DNS excluded case, the most important part of the formula is the initial grid position, followed by lap times, pit times, delta-to-lead, and completed laps. On the other hand, when DNF/DNS is included, the importance order changes to first the lap times, followed by the completed laps, initial grid position, and pit times, and finally, delta-to-lead. These insights can help teams improve their race-day strategies.
@@
TABLE I: Performance of RS and the proposed method on training and test sets in terms of accuracy (%) for $\phi$ and $\tilde{\phi}$.
Formula Method DNF/DNS excl. DNF/DNS incl.
Train Test Train Test
$\phi$ RS[10] $86.7\pm 0.6$ $82.5\pm 2.6$ $81.2\pm 0.8$ $78.8\pm 10.4$
Ours $92.9$ $81.7$ $85.3$ $85.3$
$\tilde{\phi}$ RS[10] $86.7\pm 0.7$ $82.5\pm 2.5$ $81.3\pm 0.8$ $78.9\pm 10.0$
Ours $93.9$ $81.7$ $90.2$ $64.74$
@@
Finally, we want to answer the following question: given only the laps observed thus far until lap $K$, how accurately can we predict the final standings? Figure 3 shows the evolution of the prediction accuracy of $\phi$ over the race. More specifically, when DNF/DNS are excluded, as Figure 3(a) shows, the model reaches over $85\%$ accuracy after fifteen laps. On the other hand, when DNF/DNS are included, the accuracy improves more slowly over laps, reflecting the unpredictable nature of the race events.
@@
It is worth emphasizing that the goal of this experiment is not to provide insights into Formula 1 but to serve as a proof of concept. The results highlight the abilities of temporal logic learning in analyzing and learning patterns in an interpretable and efficient manner for complex problems like learning to rank on real-world sequential data, such as racing.
@@ VI Conclusion, Limitations and Future Work
In this work, we proposed a safe learning-from-human-feedback framework, based on temporal logic formalism, that can be efficiently solved to optimality. The framework consists of two key procedures: structural pruning and $\log$-transform. Structural pruning systematically removes parts of the robustness tree that do not affect the robustness value, reducing the problem size. The $\log$-transform linearizes the robustness computation constraints with respect to weights, enabling an MILP formulation. We demonstrated the performance of the proposed method on two experiments: the preference learning problem for a robot navigation task and the learning-to-rank problem using Formula 1 data. Results show that the method can capture diverse preferences effectively.
Despite the success of the proposed method in the experiments, several factors may limit its applicability. Mainly, the method requires domain knowledge and expertise in writing temporal logic formulas to define task specifications, as well as careful hyperparameter tuning to avoid overfitting. While this need introduces a reliance on expert input, it comes with the benefit of interpretability. Looking forward, we plan to integrate more streamlined task specification methods, potentially using large language models to translate natural language descriptions to STL formulas.